# Semi-Bandit Learning for Monotone Stochastic Optimization

### v1

@@ 2.6 Framework for Continuous Distributions
Consider a stochastic problem that is up-monotone. Suppose that $\mathbf{E}=\{\mathcal{E}_{i}\}_{i=1}^{n}$ and $\mathbf{D}=\{\mathcal{D}_{i}\}_{i=1}^{n}$ are product distributions, where for each $i\in[n]$, distributions $\mathcal{D}_{i}$ and $\mathcal{E}_{i}$ satisfy:
  • $\mathcal{E}_{i}$ stochastically dominates $\mathcal{D}_{i}$.
  • For any list $\mathbf{{b}}=\langle b_{1},\cdots,b_{k-1}\rangle$ of $k-1$ thresholds, we have
    $\mathtt{TV}(\theta(\mathsf{X},\mathbf{{b}}),\theta(\mathsf{Y},\mathbf{{b}}))<\epsilon_{i},$
    where $\mathsf{X}$ (resp. $\mathsf{Y}$) is distributed according to $\mathcal{D}_{i}$ (resp. $\mathcal{E}_{i}$).
If $\sigma=\mathtt{ALG}(\mathbf{E})$ is a $k$-threshold policies and $\sigma^{*}$ is the optimal policy under $\mathbf{D}$, then:
The proof of this lemma is almost identical to that of Lemma 2.3 (see §2.2), so we just explain the changes. Recall that policy $\sigma$ corresponds to a decision tree. Using the additional fact that $\sigma$ is a $k$-threshold policy, it follows that the execution of $\sigma$ only depends on the “thresholded” r.v.s $\{\theta_{\sigma(v)}:v\in[N]\}$ where $\theta_{\sigma(v)}$ is the result of applying the thresholds (at most $k$) at node $v$ on the continuous r.v. $\mathsf{X}_{\sigma(v)}$. Crucially, these $\theta$ r.v.s are discrete and independent: so we can directly use the proof of Lemma 2.3 (where we replace the $X_{v}$ r.v.s by $\theta_{\sigma(v)}$).
Using Theorem 2.7, Lemma 2.8, and the same steps as for proving Theorem 1.1, we obtain:
Theorem 2.9.
Suppose that stochastic problem $\mathcal{P}$ has an $\alpha$-approximation algorithm via $k$-threshold policies, and $\mathcal{P}$ is either up-monotone or down-monotone. Then, there is a polynomial-time semi-bandit learning algorithm for $\mathcal{P}$ with $\alpha$-regret $O(knf_{max}\sqrt{T\log(knT)})$. Here, $n$ is the number of items, $f_{max}$ is the maximal value of the objective function and $T$ is the number of periods.
@@ 3 Applications
In this section, we show that several stochastic optimization problems are covered by our framework, resulting in $\sqrt{T\log T}$ regret online learning algorithm for all these problems.
Series Testing
@@
We start with a simple problem: there are $n$ components, where each component $i$ is “working” independently with some known probability $p_{i}$. To determine if any component $i\in[n]$ is working, we need to perform a test, which costs $c_{i}$. All $n$ components must be working for the system to be functional. The goal is to test components sequentially to determine whether/not the system is functional, at the minimum expected cost. Note that testing stops once a failed component is found: so we do not observe all the outcomes and only have semi-bandit feedback. It is easy to show that this problem is up-monotone: this is a (simple) special case of Lemma 3.4 below. It is well-known that the natural greedy policy achieves the optimal cost [But72]. So, using Theorem 1.1 with $k=2$ (all r.v.s are binary), we obtain a polynomial time online learning algorithm for series testing having $1$-regret $O(nC\sqrt{T\log T})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost.
Prophet Inequality
@@
The Prophet Inequality [KS77, SC84] is a fundamental problem in optimal stopping, which has also been used extensively in algorithmic game theory. The input consists of $n$ rewards which arrive in a given fixed sequence, say $\mathsf{X}_{1},\cdots\mathsf{X}_{n}$. Each reward $\mathsf{X}_{i}$ is drawn independently from a known distribution $\mathcal{D}_{i}$. We are interested in online policies, that upon observing each reward, selects or discards it immediately. The policy can select at most one reward, and it terminates right after making a selection (without observing any future reward). Note that we have semi-bandit feedback because only some of the rewards are observed in any policy execution. The goal is to maximize the expected selected reward. The classical results obtain a $\frac{1}{2}$-approximate policy relative to the “clairvoyant” optimal value $\mathbb{E}[\max_{i=1}^{n}\mathsf{X}_{i}]$; there are also instances where no policy can achieve a better approximation to this benchmark. Here, we will compare to a more realistic non-clairvoyant benchmark: the optimal policy which is also constrained to make selection decisions in the given order (same as an algorithm). It is known that there is an optimal threshold-based policy: given thresholds $\{\tau_{i}\}_{i=1}^{n}$, the policy selects $i$ if and only if $\mathsf{X}_{i}>\tau_{i}$. This is a $1$-threshold policy, as defined in §2.6. Moreover, the prophet inequality problem is strongly monotone (see Lemma 28 in [GHTZ21]), which implies that it is down-monotone. Using Theorem 2.9, we get a polynomial time online learning algorithm for the the prophet inequality problem with unknown distributions that has $1$-regret $O(nU\sqrt{T\log T})$ where all r.v.s are $[0,U]$ bounded. This improves upon the $O(n^{3}U\sqrt{T}\log T)$ bound from [GKSW24], although the previous result holds in the stronger bandit-feedback model. We note that there are other learning-based results [AKW14, RWW20] based on limited number of samples, that imply $\frac{1}{2}$-regret algorithms by comparing to the clairvoyant benchmark. Note that our guarantee and that of [GKSW24] are stronger because they do not incur any multiplicative approximation factor.
Combinatorial Prophet Inequalities
The basic prophet inequality concept has also been extended to settings where there is some combinatorial feasibility constraint on the selected items. Here, we have $n$ items with reward $\mathsf{X}_{i}\sim\mathcal{D}_{i}$ for each $i\in[n]$. In addition, there is a downward-closed [^4] Set family ${\cal F}$ is downward-closed if, for any $S\in{\cal F}$ we have $S^{\prime}\in{\cal F}$ for every $S^{\prime}\subseteq S$. set family ${\cal F}\subseteq 2^{[n]}$ that represents a generic feasibility constraint; the selected subset must be in ${\cal F}$. The $n$ items arrive in a given fixed sequence. When item $i$ arrives, if $S\cup\{i\}\not\in{\cal F}$ where $S$ is the set of previously selected items then item $i$ is not considered for selection (and we do not observe $\mathsf{X}_{i}$); otherwise, the policy observes the value of $\mathsf{X}_{i}$ and selects/discards item $i$. Again, note that we only have semi-bandit feedback because only a subset of items is observed by the policy. The performance of an online policy is compared to the clairvoyant optimum $\mathtt{OPT}^{*}=\mathbb{E}\left[\max_{S\in{\cal F}}\sum_{i\in S}\mathsf{X}_{i}\right]$. Many specific problems can be modeled in this manner:
  • When =${\cal F}=\{S:|S|\leq 1\}$, we get the classic prophet inequality, which has a $\frac{1}{2}$ approximation.
@@
  • When =${\cal F}=\{S:|S|\leq k\}$, there is a $1-\frac{1}{\sqrt{k+3}}$ approximation [Ala14, JMZ22].
  • When ${\cal F}$ corresponds to independent sets in a matroid, we obtain the matroid prophet inequality where again a $\frac{1}{2}$ approximation is known [KW19].
  • When ${\cal F}$ is given by the intersection of $p$ matroids, a $\frac{1}{e(p+1)}$ approximation is known [FSZ21].
  • When ${\cal F}$ is given by matchings in a graph, a $0.337$ approximation is known [EFGT22].
@@
The online policies in all these results are $1$-threshold policies (as in §2.6). For the monotone property, note that we are comparing to the clairvoyant optimum $\mathtt{OPT}^{*}$ (not the optimal policy). So, it suffices to prove down-monotonicity for $\mathtt{OPT}^{*}$, which is immediate by stochastic dominance. Therefore, using Theorem 2.9, we obtain $\alpha$-regret $O(nU\sqrt{T\log T})$ for all the above combinatorial prophet inequalities where $\alpha$ is the best (offline) approximation ratio; we assume that the r.v.s are $[0,U]$ bounded.
@@
Some approximate regret bounds can also be obtained from previous work on single-sample prophet inequalitites [AKW14, RWW20, ${}^{+}$22]. While the regret bounds via this approach are better (just $O(n)$), they need to compare to approximation ratios that are often much worse than the usual (known distribution) setting. For example, no constant-factor approximation is known for single-sample matroid prophet inequality, whereas our results imply $\frac{1}{2}$-regret of $\sqrt{T\log T}$ for general matroids, matching the approximation ratio from [KW19].
Pandora’s Box.
@@
In this problem [Wei79], we are given distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ such that r.v. $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. The realization of $\mathsf{X}_{i}$ can be ascertained by paying a known inspection cost $c_{i}$. Now, the goal is to find a policy to (adaptively) inspect a subset $S\subseteq[n]$ of the random variables to maximize $\mathbb{E}\left[\max_{i\in S}\mathsf{X}_{i}-\sum_{i\in S}c_{i}\right]$. Note that any policy only inspects a subset of items and we only receive feedback from these items, which corresponds to semi-bandit feedback. [Wei79] obtained an optimal policy based on the “reservation value” for each item and probing items according to this value until the reward for an item exceeds all remaining reservation values. The reservation value $r_{i}$ for an item $i$ is such that $\mathbb{E}[(\mathsf{X}_{i}-r_{i})_{+}]=c_{i}$. We note that this optimal policy is $1$-threshold, according to the definition in §2.6.
@@
For the online setting, we assume that the r.v.s $\mathsf{X}_{i}$ are $[0,U]$ bounded; the distributions may be discrete or continuous. It was shown in [GHTZ21] (see Lemma 31 in that paper) that the Pandora’s box problem is strongly monotone, which implies that it is down-monotone. Combined with Theorem 2.9, we get a polynomial time online learning algorithm for the Pandora’s box problem with unknown distributions that has $1$-regret $O(n(C+U)\sqrt{T\log T})$ where $C=\sum_{i\in[n]}c_{i}$ is the total cost. Our regret bound improves upon the $O(n^{4.5}(C+U)\sqrt{T}\log T)$ bound from [GKSW24], although the previous result holds in the stronger bandit-feedback model.
@@
In Pandora’s box with order constraints, in addition to the $n$ r.v.s, there are precedence constraints that enforce that any r.v. $\mathsf{X}_{i}$ may be selected only after all its predecessors have been selected. Although the original policy [Wei79] does not apply to this extension, [BFLL23] obtained a different optimal policy when the precedence constraints form a directed tree: this policy is also a 1-threshold policy. Hence, Theorem 2.9 implies an online learning algorithm for Pandora’s box with tree order constraints having $1$-regret $O(n(C+U)\sqrt{T\log T})$, where $C=\sum_{i\in[n]}c_{i}$ is the total cost and $U$ is the bound on the r.v.s.
In the matroid Pandora’s box problem [KWW16, Sin18], in addition to the $n$ r.v.s $\{\mathsf{X}_{i}\}_{i=1}^{n}$, there is a matroid $\cal M$ with groundset $[n]$. The goal is to inspect a subset $S\subseteq[n]$ of r.v.s and select a subset $B\subseteq S$ such that $B$ is independent in matroid $\cal M$. The objective is to maximize $\mathbb{E}\left[\sum_{j\in B}\mathsf{X}_{j}-\sum_{i\in S}c_{i}\right]$, the difference between the total selected value and inspection cost. We recover the original Pandora’s box problem when $\cal M$ is a rank-1 uniform matroid. There is an optimal 1-threshold policy known for this variant [Sin18]. This policy is also non-adaptive. So, by Theorem 2.9, we obtain an online learning algorithm for matroid Pandora’s box having $1$-regret $O(n(C+U)\sqrt{T\log T})$, where $C=\sum_{i\in[n]}c_{i}$ is the total cost and $U$ is the bound on the r.v.s.
Stochastic Knapsack
This is a classic problem in stochastic optimization, which was introduced in [DGV08] and has been studied extensively [BGK11, GKMR11, Ma18]. There are $n$ items with deterministic rewards $\{r_{i}\}_{i=1}^{n}$ and random costs $\{C_{i}\sim\mathcal{D}_{i}\}_{i=1}^{n}$. The realized cost $C_{i}$ of item $i$ is only known after selecting it. Given a knapsack budget $B$, a policy selects items sequentially until the total cost exceeds $B$. The objective is to maximize the expected total reward from items that fit in the knapsack. (If there is an item that overflows the budget then it does not contribute to the objective.) Note that only some subset of items is selected by a policy, and we only observe those costs as feedback. In Lemma 3.1 below, we show that this problem is down-monotone (assuming that the costs are discrete r.v.s).
There is a non-adaptive $\frac{1}{4}$-approximation algorithm for stochastic knapsack [DGV08], which probes items in a fixed sequence until the budget is exhausted; so it is a 1-threshold policy. Theorem 2.9 then implies a polynomial time online learning algorithm for stochastic knapsack having $\frac{1}{4}$-regret of $O(nR\sqrt{T\log T})$ where $R=\sum_{i=1}^{n}r_{i}$ is the total reward. There is also a better adaptive algorithm for stochastic knapsack, which is a $(\frac{1}{2}-\epsilon)$ approximation (for any constant $\epsilon>0$) [Ma18]. Theorem 1.1 then implies an online learning algorithm with $(\frac{1}{2}-\epsilon)$-regret of $O(nkR\sqrt{T\log(kT)})$ where $k$ is the maximum support size.
Our results also apply to the more general stochastic orienteering problem, where items are located at vertices in a metric space and we want to find a path with budget $B$ on the total distance (from the edges in the path) plus cost (of the visited items). There is a non-adaptive $\Omega(\frac{1}{\log\log B})$ approximation algorithm for this problem [GKNR15], which is a 1-threshold policy. Here, we obtain $\Omega(\frac{1}{\log\log B})$-regret of $O(nR\sqrt{T\log T})$.

### v2

@@ 4.3 Overall Regret
With probability at least $1-\frac{1}{nT}$, we have $\mathcal{D}_{i}\preceq_{\mathsf{SD}}\mathcal{E}^{t}_{i}$ and the total variation distance $\mathtt{TV}(\mathcal{E}_{c,i}^{t},\mathcal{D}_{c,i})\leq k\sqrt{\frac{6\log(knT)}{N_{c,i}^{t}}}$ for $c\in[k]$, $i\in[n]$, and $t\in[T]$.
Proof.
For any $c\in[k]$, $i\in[n]$, and $t\in[T]$, let $B_{c,i}^{t}$ denote the event that the stated condition fails to hold for the triple $(c,i,t)$. By Theorem 4.3 with $\delta=\frac{2}{k^{2}n^{3}T^{3}}$, we have:
$\Pr\left(B_{c,i}^{t}\right)\leq\sum_{m=1}^{T}\Pr\left(B_{c,i}^{t}\wedge N_{c,i}^{t}=m\right)\leq\delta T.$
Applying another union bound over all values of $c$, $i$, and $t$, we obtain:
$\Pr\left(\vee_{c=1}^{k}\vee_{i=1}^{n}\vee_{t=1}^{T}B_{c,i}^{t}\right)\leq\sum_{c=1}^{k}\sum_{i=1}^{n}\sum_{t=1}^{T}\Pr\left(B_{c,i}^{t}\right)\leq knT^{2}\delta\leq\frac{1}{nT},$
where the last inequality uses the assumption that $n\geq 2$. ∎
We now define a good event $G$ that corresponds to the condition in Lemma 4.10 holding true for all $c\in[k]$, $i\in[n]$ and $t\in[T]$. Next, we complete the proof assuming that $G$ holds.
Bounding the regret as a sum over time $t$.
To bound the overall regret, it suffices to bound the expected regret $R^{t}=f(\sigma^{t})-\alpha\cdot f(\sigma^{*})$ at each time $t\in[T]$. For each time $t\in[T]$, using the good event, we can apply Lemma 4.8 with distributions $\mathbf{E}^{t}$, $\mathbf{D}$, and parameters $\epsilon_{c,i}^{t}=k\cdot\sqrt{\frac{6\log(knT)}{N_{c,i}^{t}}}$ to obtain
$R^{t}\leq f_{\max}\sum_{i=1}^{n}\sum_{c=1}^{k}{Q}_{c,i}(\sigma^{t})\cdot\epsilon_{c,i}^{t}.$ (34)
So, the overall regret is
$\alpha\text{-}R(T)\leq f_{\max}\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}\sum_{c=1}^{k}{Q}_{c,i}(\sigma^{t})\cdot\epsilon_{c,i}^{t}\right]=kf_{\max}\sqrt{6\log(knT)}\cdot\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}\sum_{c=1}^{k}\frac{{Q}_{c,i}(\sigma^{t})}{\sqrt{N^{t}_{c,i}}}\right].$
Recall that $\mathbf{h}^{t-1}=(\mathbf{x}^{1},\cdots\mathbf{x}^{t-1})$ is the history until time $t$. It now suffices to show
$\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\sum_{i=1}^{n}\sum_{c=1}^{k}\frac{{Q}_{c,i}(\sigma^{t})}{\sqrt{N^{t}_{c,i}}}\right]\leq 2n\sqrt{kT}.$ (35)
Proving (6) as a sum over decision paths.
Similar to the semi-bandit case, we define:
[…]
Above, $\mathbf{I}$ is the indicator function. By linearity of expectation, we have
[…]
$=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{1}{\sqrt{N^{t}_{c,i}}}\cdot\Pr_{\mathbf{x}^{t}}[i\text{ probed by }\sigma^{t}\text{ with }a_{c}]\right]\,\,=\,\,\sum_{t=1}^{T}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{Q_{c,i}(\sigma^{t})}{\sqrt{N^{t}_{c,i}}}\right],$ (36)
where the second equality uses the fact that event {$i$ probed by $\sigma^{t}$ with cc} only depends on $\mathbf{h}^{t}=(\mathbf{h}^{t-1},\mathbf{x}^{t})$, the third equality uses the fact that $N^{t}_{c,i}$ only depends on $\mathbf{h}^{t-1}$ and that $\mathbf{x}^{t}$ is independent of $\mathbf{h}^{t-1}$, and the last equality is by the definition of $Q_{c,i}(\sigma^{t})$ and the fact that $\mathbf{x}^{t}\sim\mathbf{D}$.
Using (36) and adding over $i\in[n]$ and $c\in[k]$, we get
$\sum_{t=1}^{T}\sum_{c=1}^{k}\sum_{i=1}^{n}\mathbb{E}_{\mathbf{h}^{t-1}}\left[\frac{{Q}_{c,i}(\sigma^{t})}{\sqrt{N^{t}_{i}}}\right]=\sum_{i=1}^{n}\sum_{c=1}^{k}\mathbb{E}_{\mathbf{h}^{T}}\left[Z_{i}^{c}(\mathbf{h}^{T})\right].$
Therefore, proving (35) is equivalent to:
$\sum_{i=1}^{n}\sum_{c=1}^{k}\mathbb{E}_{\mathbf{h}^{T}}[Z_{i}^{c}(\mathbf{h}^{T})]\leq 2n\sqrt{kT}.$ (37)
Now, for any $i\in[n]$ and $c\in[k]$, we obtain:
$Z_{i}^{c}(\mathbf{h}^{T})=\sum_{t=1}^{T}\frac{\mathbf{I}\left[i\text{ probed by }\sigma^{t}(\mathbf{h}^{t-1})\text{ with }a_{c}\right]}{\sqrt{N^{t}_{c,i}(\mathbf{h}^{t-1})}}\leq\sum_{t=1}^{N^{T}_{c,i}}\frac{1}{\sqrt{t}}\leq 2\sqrt{N^{T}_{c,i}}.$
The first inequality uses the fact that if item $i$ gets probed with threshold $a_{c}$ at time $t$ then $N_{c,i}^{t}=N_{c,i}^{t-1}+1$. Using Jensen’s inequality and the fact that the total number of probes for one item $i$ is at most $T$, we have:
$\sum_{c=1}^{k}\sqrt{N^{T}_{c,i}}\leq k\sqrt{\frac{\sum_{c=1}^{k}N^{T}_{c,i}}{k}}\leq\sqrt{kT}.$
This completes the proof of (37) and hence (35).
We note that our good event $G$ holds with probability at least $1-\frac{1}{nT}$. As in the semi-bandit setting, we can convert the high-probability regret bound into an expected regret bound. This concludes the proof of Theorem 4.1.
The proof of Theorem 4.2 is identical: we just need to use the appropriate sampling result (Theorem 4.7) and stability lemma (Lemma 4.9).
@@ 5 Applications
In this section, we show that several stochastic optimization problems are covered by our framework, resulting in $\sqrt{T\log T}$ regret online learning algorithm for all these problems. In each of these problems, the distributions of the random parameters are unknown to the online algorithm; all other (deterministic) parameters are known. We summarize the overall guarantees in Table 1, highlighting the dependence on $n$, $T$, the support size $k$ of the underlying distributions (where applicable), and the distributional assumptions needed. See the corresponding subsection for more details and exact bounds.
Application Approx. Our $\alpha$-Regret Distributional
Factor ($\alpha$) Bound Assumptions
Series Testing 11 $O(n\sqrt{T\log T})$ Bernoulli
Prophet Inequality 11 $O(n\sqrt{T\log T})$ continuous
Multi-Item Prophet Inequality $1-\frac{1}{\sqrt{\kappa+3}}$
Matroid Prophet Inequality $1/2$
Pandora’s Box 11 $O(n\sqrt{T\log T})$ continuous
Stochastic Knapsack $1/2-\epsilon$ $O(nk\sqrt{T\log T})$ discrete
Stochastic Orienteering $\Omega(1/\log\log B)$ $O(nk\sqrt{T\log T})$ discrete
Unweighted Stochastic Matching $1/2$ $O(n\sqrt{T\log T})$ Bernoulli
Weighted Stochastic Matching $0.382$
Stochastic Covering Knapsack 33 $O(nk\sqrt{T\log T})$ discrete
Stochastic $\kappa$-TSP $O(1)$ $O(nk\sqrt{T\log T})$ discrete
Stochastic Submodular Maximization $1-\frac{1}{e}$ $O(n\sqrt{T\log T})$ Bernoulli
Stochastic Submodular Cover $1+\ln(Q)$
Single Resource Revenue Management 11 $O(nk^{2}\sqrt{kT\log(kT)})$ discrete
FSPM: Cardinality Constrained 11 $O(n\sqrt{kT\log(kT)})$ discrete
ASPM: Cardinality Constrained $1-\epsilon$
ASPM: Matroid Constrained $1-\frac{1}{e}$
Table 1: Summary of Regret Bounds; Fixed-Order Sequential Posted-Pricing Mechanisms,
Adaptive Sequential Posted-Pricing Mechanisms
5.1 Series Testing
@@
We start with a simple problem: there are $n$ components, where each component $i$ is “working” independently with some known probability $p_{i}$. To determine if any component $i\in[n]$ is working, we need to perform a test, which costs $c_{i}$. All $n$ components must be working for the system to be functional. The goal is to test components sequentially to determine whether/not the system is functional, at the minimum expected cost. Note that testing stops once a failed component is found: so we do not observe all the outcomes and only have semi-bandit feedback. It is easy to show that this problem is up-monotone: this is a (simple) special case of Lemma 5.4 below. It is well-known that the natural greedy policy achieves the optimal cost (Butterworth, 1972). So, using Theorem 1.1 with $k=2$ (all r.v.s are binary), we obtain a polynomial time online learning algorithm for series testing having $1$-regret $O(nC\sqrt{T\log T})$ where $C=\sum_{i=1}^{n}c_{i}$ is the total cost.
5.2 Prophet Inequality
@@
The Prophet Inequality (Krengel and Sucheston, 1977; Samuel-Cahn, 1984) is a fundamental problem in optimal stopping, which has also been used extensively in algorithmic game theory. The input consists of $n$ rewards which arrive in a given fixed sequence, say $\mathsf{X}_{1},\ldots,\mathsf{X}_{n}$. Each reward $\mathsf{X}_{i}$ is drawn independently from a known distribution $\mathcal{D}_{i}$. We are interested in online policies, that upon observing each reward, selects or discards it immediately. The policy can select at most one reward, and it terminates right after making a selection (without observing any future reward). Note that we have semi-bandit feedback because only some of the rewards are observed in any policy execution. The goal is to maximize the expected selected reward. The classical results obtain a $\frac{1}{2}$-approximate policy relative to the “clairvoyant” optimal value $\mathbb{E}[\max_{i=1}^{n}\mathsf{X}_{i}]$; there are also instances where no policy can achieve a better approximation to this benchmark. Here, we will compare to a more realistic non-clairvoyant benchmark: the optimal policy which is also constrained to make selection decisions in the given order (same as an algorithm). It is known that there is an optimal threshold-based policy: given thresholds $\{\tau_{i}\}_{i=1}^{n}$, the policy selects $i$ if and only if $\mathsf{X}_{i}>\tau_{i}$. This is a $1$-threshold policy, as defined in Section 3. Moreover, the prophet inequality problem is strongly monotone (see Lemma 28 in Guo et al. (2021)), which implies that it is down-monotone. Finally, for any policy $\sigma$ and leaf $\ell$ (see (23) for details), the corresponding reward function is $f_{\sigma,\ell}(\mathbf{x})=x_{\ell},$ where $x_{\ell}$ denotes the value of the random variable at leaf $\ell$. It is straightforward to verify that $f_{\sigma,\ell}$ is coordinate-wise monotone (see Remark 3.8). Using Theorem 3.7, we get a polynomial time online learning algorithm for the the prophet inequality problem with unknown distributions that has $1$-regret $O(nU\sqrt{T\log T})$ where all r.v.s are $[0,U]$ bounded. This improves upon the $O(n^{3}U\sqrt{T}\log T)$ bound from Gatmiry et al. (2024), although the previous result holds in the stronger bandit-feedback model. We note that there are other learning-based results (Azar et al., 2014; Rubinstein et al., 2020) based on limited number of samples, that imply $\frac{1}{2}$-regret algorithms by comparing to the clairvoyant benchmark. Note that our guarantee and that of Gatmiry et al. (2024) are stronger because they do not incur any multiplicative approximation factor.
Combinatorial Prophet Inequalities.
[…]
  • When ${\cal F}=\{S:|S|\leq 1\}$, we get the classic prophet inequality, which has a $\frac{1}{2}$ approximation.
@@
  • When ${\cal F}=\{S:|S|\leq\kappa\}$, there is a $1-\frac{1}{\sqrt{\kappa+3}}$ approximation (Alaei, 2014; Jiang et al., 2022).
[…]
@@
For the monotone property, note that we are comparing to the clairvoyant optimum $\mathtt{OPT}^{*}$ (not the optimal policy). So, it suffices to prove down-monotonicity for $\mathtt{OPT}^{*}$, which is immediate by stochastic dominance. The online policies in all these results are $1$-threshold policies. For any fixed policy $\sigma$ and leaf $\ell$ (see (23)), we have
$f_{\sigma,\ell}(\mathbf{x})=\sum_{i=1}^{n}x_{i}\cdot\mathbf{I}_{i\in{\cal F}_{\ell}},$ (38)
[…]
@@
Some approximate regret bounds can also be obtained from previous work on single-sample prophet inequalities (Azar et al., 2014; Rubinstein et al., 2020; Caramanis et al., 2022). While the regret bounds via this approach are better (independent of $T$), they need to compare to approximation ratios that are often worse than the usual (known distribution) setting. In particular, Fu et al. (2024) obtained an approximation ratio of $(\frac{1}{4}-\epsilon)$ for general matroids using $O_{\epsilon}(n\log^{4}n)$ total samples, which implies a learning algorithm having $\left(\frac{1}{4}-\epsilon\right)$-regret of $O_{\epsilon}(n\log^{4}n)$. In contrast, our results imply $\frac{1}{2}$-regret of $\sqrt{T\log T}$, exactly matching the best-possible approximation ratio from Kleinberg and Weinberg (2019).
5.3 Pandora’s Box
@@
In this problem (Weitzman, 1979), we are given distributions $\mathcal{D}_{1},\ldots,\mathcal{D}_{n}$ such that r.v. $\mathsf{X}_{i}\sim\mathcal{D}_{i}$. The realization of $\mathsf{X}_{i}$ can be ascertained by paying a known inspection cost $c_{i}$. Now, the goal is to find a policy to (adaptively) inspect a subset $S\subseteq[n]$ of the random variables to maximize $\mathbb{E}\left[\max_{i\in S}\mathsf{X}_{i}-\sum_{i\in S}c_{i}\right]$. Note that any policy only inspects a subset of items and we only receive feedback from these items, which corresponds to semi-bandit feedback. Weitzman (1979) obtained an optimal policy based on the “reservation value” for each item and probing items according to this value until the reward for an item exceeds all remaining reservation values. The reservation value $r_{i}$ for an item $i$ is such that $\mathbb{E}[(\mathsf{X}_{i}-r_{i})_{+}]=c_{i}$. We note that this optimal policy is $1$-threshold.
@@
For the online setting, we assume that the r.v.s $\mathsf{X}_{i}$ are $[0,U]$ bounded; the distributions may be discrete or continuous. It was shown in Guo et al. (2021) (see Lemma 31 in that paper) that the Pandora’s box problem is strongly monotone, which implies that it is down-monotone. Similar to Prophet Inequality, for any policy $\sigma$ and leaf $\ell$ , the corresponding reward function is $f_{\sigma,\ell}(\mathbf{x})=\max_{i\in S_{\ell}}x_{i}-\sum_{i\in S_{\ell}}c_{i},$ where $S_{\ell}$ is the set of inspected r.v.s at leaf $\ell$. Hence $f_{\sigma,\ell}$ is coordinate-wise monotone. Then combined with Theorem 3.7, we get a polynomial time online learning algorithm for the Pandora’s box problem with unknown distributions that has $1$-regret $O(n(C+U)\sqrt{T\log T})$ where $C=\sum_{i\in[n]}c_{i}$ is the total cost. Our regret bound improves upon the $O(n^{4.5}(C+U)\sqrt{T}\log T)$ bound from Gatmiry et al. (2024), although the previous result holds in the stronger bandit-feedback model.
@@
In Pandora’s box with order constraints, in addition to the $n$ r.v.s, there are precedence constraints that enforce that any r.v. $\mathsf{X}_{i}$ may be selected only after all its predecessors have been selected. Although the original policy of Weitzman (1979) does not apply to this extension, Boodaghians et al. (2023) obtained a different optimal policy when the precedence constraints form a directed tree: this policy is also a 1-threshold policy. Hence, Theorem 3.7 implies an online learning algorithm for Pandora’s box with tree order constraints having $1$-regret $O(n(C+U)\sqrt{T\log T})$, where $C=\sum_{i\in[n]}c_{i}$ is the total cost and $U$ is the bound on the r.v.s.
[…]
5.4 Stochastic Knapsack
This is a classic problem in stochastic optimization, which was introduced in Dean et al. (2008) and has been studied extensively since (Bhalgat et al., 2011; Gupta et al., 2011; Ma, 2018). There are $n$ items with deterministic rewards $\{r_{i}\}_{i=1}^{n}$ and random costs $\{C_{i}\sim\mathcal{D}_{i}\}_{i=1}^{n}$. The realized cost $C_{i}$ of item $i$ is only known after selecting it. Given a knapsack budget $B$, a policy selects items sequentially until the total cost exceeds $B$. The objective is to maximize the expected total reward from items that fit in the knapsack. If there is an item that overflows the budget then it does not contribute to the objective. Note that only some subset of items is selected by a policy, and we only observe those costs as feedback. In Lemma 5.1 below, we show that this problem is up-monotone (assuming that the costs are discrete r.v.s). There is an adaptive algorithm for stochastic knapsack, which is a $(\frac{1}{2}-\epsilon)$ approximation (for any $\epsilon>0$(Ma, 2018). Theorem 1.1 then implies an online learning algorithm with $(\frac{1}{2}-\epsilon)$-regret of $O(nkR\sqrt{T\log T})$ where $k$ is the maximum support size.
Our results also apply to the more general stochastic orienteering problem, where items are located at vertices in a metric space and we want to find a path with budget $B$ on the total distance (from the edges in the path) plus cost (of the visited items). There is a non-adaptive $\Omega(\frac{1}{\log\log B})$ approximation algorithm for this problem (Gupta et al., 2015). Here, we obtain $\Omega(\frac{1}{\log\log B})$-regret of $O(nkR\sqrt{T\log T})$.